OpenStudy (anonymous):

A manufacture has been selling 1150 television sets a week at 480 dollars each. A market survey indicates that for each 12 -dollar rebate offered to a buyer, the number of sets sold will increase by 120 per week. a) Find the demand function p(x), where x is the number of the television sets sold per week, assuming that p(x) is linear. p(x)= b) How large of a rebate should the company offer to a buyer, in order to maximize its revenue? c) If the weekly cost function is 92000 + 160 x, how should it set the size of the rebate to maximize its profit?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

help any one

OpenStudy (anonymous):

P(x) = - x/10 + 115 + 480 = - x/10 + 595

OpenStudy (anonymous):

why x/10?

OpenStudy (anonymous):

slope = -12/120

OpenStudy (anonymous):

ok so what would be the demand func?

OpenStudy (anonymous):

P(x) = = - x/10 + 595

OpenStudy (anonymous):

Now can you find Revenue func?

OpenStudy (anonymous):

im trying

OpenStudy (anonymous):

can u please explain how u got p(x)

OpenStudy (anonymous):

x = quantity of TV set, y = unit price => (x, y ) = ( 1,150, 480 )

OpenStudy (anonymous):

y - 480 = (- 1/10 ) ( x - 1,150 )

OpenStudy (anonymous):

Can you do the calculation?

OpenStudy (anonymous):

are u finding the eq of the line?

OpenStudy (anonymous):

"assuming that p(x) is linear"

OpenStudy (anonymous):

ok now i understood the whole equation of the line...

OpenStudy (anonymous):

so how would i solve part b?

OpenStudy (anonymous):

R(x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would R(X) be P(X)*x?

OpenStudy (anonymous):

You just has responded to me in the very previous post like that :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i suck at math...

OpenStudy (anonymous):

At first you need to master the business concept, otherwise i give up :(

OpenStudy (anonymous):

ok so this what i got for R(x)=(-x*(x-5950))/(10)

OpenStudy (anonymous):

Multiply it out, will you!

OpenStudy (anonymous):

(-x^2+5950)/(10)

OpenStudy (anonymous):

R(x) = -x²/ 10 + 595

OpenStudy (anonymous):

What the problem want you to find?

OpenStudy (anonymous):

How large of a rebate should the company offer to a buyer, in order to maximize its revenue?

OpenStudy (anonymous):

R(x) = -x²/ 10 + 595 should = max revenue

OpenStudy (anonymous):

How to find max?

OpenStudy (anonymous):

R(x) = -x²/ 10 + 595 is the rq for max R

OpenStudy (anonymous):

eq*

OpenStudy (anonymous):

R'(x) = ...

OpenStudy (anonymous):

R(x) = -x²/ 10 + 595x

OpenStudy (anonymous):

any ideas on how to calc the rebate?

OpenStudy (anonymous):

They says find R'(x)

OpenStudy (anonymous):

R'(x) = 595-(x/5)

OpenStudy (anonymous):

Set R'(x) = 0 ==> x = ..

OpenStudy (anonymous):

x= 2975

OpenStudy (anonymous):

but this is not the amt for reabte ? so what now ?

OpenStudy (anonymous):

$12 rebate --> 120 set TV

OpenStudy (anonymous):

x= 2975 is the number of tv sets solf to max revenue?

OpenStudy (anonymous):

I think multiply by 1/10 ~ $298 rebate!

OpenStudy (anonymous):

no thats not it :(

OpenStudy (anonymous):

What the result?

OpenStudy (anonymous):

i dont know what the result is ::(

OpenStudy (anonymous):

So how do you know it's not $298?

OpenStudy (anonymous):

ok if 12rebate = 120 tv sets and x= 2975, then divide 2975/120 and mulitply by 12

OpenStudy (anonymous):

this gives us 297.5 or 298 and my hw says this is wrong some how

OpenStudy (anonymous):

my hw is online btw

OpenStudy (anonymous):

Can you post the sample, so I can have the idea?

OpenStudy (anonymous):

A manufacture has been selling 1150 television sets a week at 480 dollars each. A market survey indicates that for each 12 -dollar rebate offered to a buyer, the number of sets sold will increase by 120 per week. a) Find the demand function p(x), where x is the number of the television sets sold per week, assuming that p(x) is linear. p(x)= b) How large of a rebate should the company offer to a buyer, in order to maximize its revenue? c) If the weekly cost function is 92000 + 160 x, how should it set the size of the rebate to maximize its profit?

OpenStudy (anonymous):

thats the entire problem as it is from the website

OpenStudy (anonymous):

i dont have any other sample...sorry

OpenStudy (anonymous):

You should learn the debate topic before having the homework, right?

OpenStudy (anonymous):

yes but this problem is tricky

OpenStudy (anonymous):

ok i got the rebate ...its 182.5/ after trying several attempts to solve it ..

OpenStudy (anonymous):

Let me think !

OpenStudy (anonymous):

Yep, you're right :D

OpenStudy (anonymous):

ok great team work

OpenStudy (anonymous):

now can u help me with part c???

OpenStudy (anonymous):

Hey, you're not suck at math! I believe so ;)

OpenStudy (anonymous):

lol ok thanks...but u have been a great help too

OpenStudy (anonymous):

The proof is whatever I ask, you can respond smoothly :)

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

Now, what is Profit?

OpenStudy (anonymous):

profit = revenue - cost

OpenStudy (anonymous):

Put them into the func, will you!

OpenStudy (anonymous):

Trust me, if you're suck at math, you already run away from me :P

OpenStudy (anonymous):

ok a function would be ((x^2/10)+595x)-(92000+160x)

OpenStudy (anonymous):

this calculates to be (x^2/10)+775x-92000

OpenStudy (anonymous):

Then maximize profit means ...

OpenStudy (anonymous):

Actually, P(x) = - x²/ 10 + 435x - 9,200

OpenStudy (anonymous):

Hey, where ARE you?

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

just trying to figure it out

OpenStudy (anonymous):

Just take derivative P'(x)

OpenStudy (anonymous):

ok lets see

OpenStudy (anonymous):

= 755-(x/5)

OpenStudy (anonymous):

P(x) = - x²/ 10 + 435x - 9,200

OpenStudy (anonymous):

= 755-(x/5) is the derivative

OpenStudy (anonymous):

Your calculation of Profit func is incorrect!

OpenStudy (anonymous):

o no

OpenStudy (anonymous):

Just fix the subtraction 595x - 160x = 435x

OpenStudy (anonymous):

x/5+755 ...is this correct?

OpenStudy (anonymous):

Now take derivative Profit func P(x) = - x²/ 10 + 435x - 9,200 P '(x) = ...

OpenStudy (anonymous):

Exactly what you did for the last b) question!

OpenStudy (anonymous):

ok derivative is x/5 +435

OpenStudy (anonymous):

x=-2175

OpenStudy (anonymous):

P'(x) = - x/5 + 435 = 0 => x = 2,175 units

OpenStudy (anonymous):

Now what have you done to get the rebate?

OpenStudy (anonymous):

ok 102.5

OpenStudy (anonymous):

thats it/ thanks so much for working with me to solve this...

OpenStudy (anonymous):

I got $262,5 !

OpenStudy (anonymous):

no rebate is 102.5

OpenStudy (anonymous):

How do you calculate?